Hi,
I have a maths question where the hypotenuse is 10m but the other to sides are unknown. These sides are both equal though and i have to answer in 2dp.

Answer by nerdybill(7384)

Let x = length of one side
then from Pythagorean theorem:
x^2 + x^2 = 10^2
2x^2 = 100
x^2 = 50
x = sqrt(50)
x = sqrt(5*5*2)
x = 5*sqrt(2)
x = 7.07 meters